# arXiv:2603.02473v1[cs.AI] 02 Mar 2026

### v1

@@ Abstract
Memory-augmented LLM agents store and retrieve information from prior interactions, yet the relative importance of how memories are written versus how they are retrieved remains unclear. We introduce a diagnostic framework that analyzes how performance differences manifest across write strategies, retrieval methods, and memory utilization behavior, and apply it to a $3\times 3$ study crossing three write strategies (raw chunks, Mem0-style fact extraction, MemGPT-style summarization) with three retrieval methods (cosine, BM25, hybrid reranking). On LoCoMo, retrieval method is the dominant factor: average accuracy spans 20 points across retrieval methods (57.1% to 77.2%) but only 3–8 points across write strategies. Raw chunked storage, which requires zero LLM calls, matches or outperforms expensive lossy alternatives, suggesting that current memory pipelines may discard useful context that downstream retrieval mechanisms fail to compensate for. Failure analysis shows that performance breakdowns most often manifest at the retrieval stage rather than at utilization. We argue that, under current retrieval practices, improving retrieval quality yields larger gains than increasing write-time sophistication. Code is publicly available at https://github.com/boqiny/memory-probe.
@@ 1 Introduction
A growing body of work equips LLM agents with persistent memory [6, 9, 1]. These systems differ mainly in what they store: some keep raw conversation text [4], others extract structured facts with conflict resolution [1, 9], and still others compress sessions into summaries [6]. Recent work has pushed further with inter-memory linking [9] and reinforcement learning for end-to-end memory management [11, 7]. Yet a basic question remains open: does the write strategy actually matter for downstream performance, or do performance differences primarily arise at retrieval? Current benchmarks [5, 8] only measure end-to-end accuracy, making it hard to tell whether errors stem from what was stored, how it was retrieved, or how the LLM used the context (Figure 2). We review related work in Appendix A.
@@ 2.1 Write Strategies
We evaluate three prompt-based write strategies on LoCoMo [5] (10 multi-session conversations, $\sim$600 turns each, $\sim$200 questions per conversation). All configurations use GPT-5-mini as the backbone LLM and text-embedding-3-small (1536-d) for embeddings.
Basic RAG stores raw 3-turn conversation chunks with speaker names and timestamps, requiring zero LLM calls at write time. Extracted Facts follows Mem0 [1]: an LLM extracts self-contained facts per session, followed by embedding-based matching and LLM conflict resolution (add/update/noop). Summarized Episodes follows MemGPT [6]: each session is compressed into a single summary paragraph. Table 3 summarizes computational costs; full prompts and implementation details are provided in Appendix C.
@@ 2.2 Retrieval Methods
We pair each write strategy with three retrieval methods of increasing sophistication, with a default retrieval budget of k=5k{=}5. Cosine similarity returns the top-$k$ memory entries by cosine distance between the query embedding and each stored entry’s embedding; this is the default method in most deployed memory systems [1, 9] and captures semantic similarity, but may miss lexically relevant entries with different surface forms. BM25 scores entries by term-frequency-based keyword overlap, complementing embedding-based retrieval by surfacing entries that share exact terms with the query; however, it fails when relevant memories use different vocabulary than the question. Hybrid+Rerank first pools the top-$2k$ candidates from both cosine and BM25, then uses GPT-5.2 as an LLM judge to rerank the union down to top-$k$; this adds one LLM call per query but leverages both semantic and lexical signals, with the reranker resolving disagreements between the two retrieval sources.
@@ Appendix A Related Work
Memory systems for LLM agents span a design spectrum from raw storage to learned management. MemGPT [6] uses recursive summarization; Mem0 [1] extracts facts with conflict resolution; A-MEM [9] adds inter-memory linking; and MemoryBank [10] combines raw logs with hierarchical summaries and forgetting curves. MEM1 [11] and Mem-$\alpha$ [7] learn memory management via reinforcement learning, falling outside our prompt-based scope. Several benchmarks evaluate long-term memory, including LoCoMo [5], LongMemEval [8], MemoryAgentBench [2], and RealTalk [3], but all measure downstream accuracy without decomposing why a system fails. Our probing framework fills this gap.

### v2

# Diagnosing Retrieval vs. Utilization Bottlenecks in LLM Agent Memory

@@ Abstract
Memory-augmented LLM agents store and retrieve information from prior interactions, yet the relative importance of how memories are written versus how they are retrieved remains unclear. We introduce a diagnostic framework that analyzes how performance differences manifest across write strategies, retrieval methods, and memory utilization behavior, and apply it to a $3\times 3$ study crossing three write strategies (raw chunks, Mem0-style fact extraction, MemGPT-style summarization) with three retrieval methods (cosine, BM25, hybrid reranking). On LoCoMo, retrieval method is the dominant factor: average accuracy spans 20 points across retrieval methods (57.1% to 77.2%) but only 3–8 points across write strategies. Raw chunked storage, which requires zero LLM calls, matches or outperforms expensive lossy alternatives, suggesting that current memory pipelines may discard useful context that downstream retrieval mechanisms fail to compensate for. Failure analysis shows that performance breakdowns most often manifest at the retrieval stage rather than at utilization. We argue that, under current retrieval practices, improving retrieval quality yields larger gains than increasing write-time sophistication.
@@ 1 Introduction
A growing body of work equips LLM agents with persistent memory (Packer et al., 2023; Xu et al., 2025; Chhikara et al., 2025). These systems differ mainly in what they store: some keep raw conversation text (Lewis et al., 2020), others extract structured facts with conflict resolution (Chhikara et al., 2025; Xu et al., 2025), and still others compress sessions into summaries (Packer et al., 2023). Recent work has pushed further with inter-memory linking (Xu et al., 2025) and reinforcement learning for end-to-end memory management (Zhou et al., 2025; Wang et al., 2025). Yet a basic question remains open: does the write strategy actually matter for downstream performance, or do performance differences primarily arise at retrieval? Current benchmarks (Maharana et al., 2024; Wu et al., 2025) only measure end-to-end accuracy, making it hard to tell whether errors stem from what was stored, how it was retrieved, or how the LLM used the context (Figure 2). We review related work in Appendix A.
@@ 2.1 Write Strategies
We evaluate three prompt-based write strategies on LoCoMo (Maharana et al., 2024) (10 multi-session conversations, $\sim$600 turns each, $\sim$200 questions per conversation). All configurations use GPT-5-mini as the backbone LLM and text-embedding-3-small (1536-d) for embeddings.
Basic RAG stores raw 3-turn conversation chunks with speaker names and timestamps, requiring zero LLM calls at write time. Extracted Facts follows Mem0 (Chhikara et al., 2025): an LLM extracts self-contained facts per session, followed by embedding-based matching and LLM conflict resolution (add/update/noop). Summarized Episodes follows MemGPT (Packer et al., 2023): each session is compressed into a single summary paragraph. Table 3 summarizes computational costs; full prompts and implementation details are provided in Appendix C.
@@ 2.2 Retrieval Methods
We pair each write strategy with three retrieval methods of increasing sophistication, with a default retrieval budget of k=5k{=}5. Cosine similarity returns the top-$k$ memory entries by cosine distance between the query embedding and each stored entry’s embedding; this is the default method in most deployed memory systems (Chhikara et al., 2025; Xu et al., 2025) and captures semantic similarity, but may miss lexically relevant entries with different surface forms. BM25 scores entries by term-frequency-based keyword overlap, complementing embedding-based retrieval by surfacing entries that share exact terms with the query; however, it fails when relevant memories use different vocabulary than the question. Hybrid+Rerank first pools the top-$2k$ candidates from both cosine and BM25, then uses GPT-5.2 as an LLM judge to rerank the union down to top-$k$; this adds one LLM call per query but leverages both semantic and lexical signals, with the reranker resolving disagreements between the two retrieval sources.
@@ Appendix A Related Work
Memory systems for LLM agents span a design spectrum from raw storage to learned management. MemGPT (Packer et al., 2023) uses recursive summarization; Mem0 (Chhikara et al., 2025) extracts facts with conflict resolution; A-MEM (Xu et al., 2025) adds inter-memory linking; and MemoryBank (Zhong et al., 2023) combines raw logs with hierarchical summaries and forgetting curves. MEM1 (Zhou et al., 2025) and Mem-$\alpha$ (Wang et al., 2025) learn memory management via reinforcement learning, falling outside our prompt-based scope. Several benchmarks evaluate long-term memory, including LoCoMo (Maharana et al., 2024), LongMemEval (Wu et al., 2025), MemoryAgentBench (Hu et al., 2025), and RealTalk (Lee et al., 2025), but all measure downstream accuracy without decomposing why a system fails. Our probing framework fills this gap.